log(2x)=2logx I believe this one is true. Am I right? Thank you.
Answer by venugopalramana(3286)

log(2x)=2logx I believe this one is true. Am I right? Thank you.
NO THE FORMULA IS LOG(A*B)=LOG(A)+LOG(B)
SO......LOG(2X)=LOG(2)+LOG(X)